HI!
I just started Algebra 2 this year, and we are reviewing from algebra right now, and I can not remember how to solve a problem where I am isolating a variable. The problem is:
x= 2y^2/5 +1
I am trying to get y alone.
the 2y^2 is all over 5, so its like a fraction. So far, I have put the 1 over with the x, so now it looks like
-1+x = 2y^2/5
Then I multipled the right side by 5, making that drop off, so now it looks like:
-1x= 2y^2. Now, I am stuck...How do I solve from here?

Answer by stanbon(48577)

-1x= 2y^2.
(-1/2)x = y^2
y = sqrt[(-1/2)x] or y = -sqrt[(-1/2)x]
----------------------------------------------
Comment:
I doubt this is the answer you are looking for, but it is
the answer to the problem you have described.
----------------------------------------------
Cheers,
Stan H.